OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 3y = 7 3x - 5y = -23 -1 1 -4 4

OpenStudy (anonymous):

HELP!!!!!!

OpenStudy (anonymous):

sorry i dont know but i think it -4

OpenStudy (anonymous):

thanks anyways

OpenStudy (shane_b):

You can solve this a couple of ways. I'll use the substitution method: Original equations:\[5x+3y=7\]\[3x-5y=-23\] Solve one of those equations for a single variable. Since the question is only asking for the y value, I'll choose to solve the first equation for x:\[x=\frac{7-3y}{5}\]Substitute that into the second equation anywhere you see an x:\[3(\frac{7-3y}{5})-5y=-23\]Distribute the -3 to the fraction and you get:\[\frac{21-9y}{5}-5y=-23\] Solve for y to get y = 4.